An enzyme catalyzes a reaction with a Km of 8.50 mM and a Vmax of 2.70 mM 5-1. Calculate the reaction velocity, vo, for each[S] = 13.0 mM Vo: 2.05 mm. s-1
Answer #1

for enzyme kinetics,

Vo = Vmax. [S] /(Km + [S]). (1)

given Km = 8.50mM.  

Vmax =2.70mM/s

(i) for [S] = 2.25mM

put values in (1)

Vo = 2.70 * 2.25/(8.50 + 2.25) = 0.565mM/s = 0.57mM/s

(ii) for [S] = 8.50mM

put values in (1)

Vo = 2.70 *8.50/(8.50+ 8.50) = 1.35 mM/s

(iii) for [S] =13mM

put values in (1)

Vo = 2.70 * 13/(8.50+13) = 1.63

hence we calculated Vo for different values of [S].